# Some statistical aspects of the Covid-19 response

### v2

@@ 4.1 Person-to-person variability in SEIR models
[…]
@@
[…]
@@
[…]
@@ 4.2 Other major modelling omissions
One such consequence concerns the modelling of $R$ after lockdown. $R$ measures the average number of new infections caused by each existing infection. Crucially it is the population of infections that is being averaged over, not the population of people. Immediately after lockdown $R$ is depressed in the locked-down population, increasingly so with time as the infectious run out of household members to infect. Meanwhile the key-worker population maintains a higher $R$. Since initially most infections are in the locked-down compartment the whole population $R$ initially tracks what is happening in that compartment, but over time an ever increasing proportion of infections are in the key-worker population, so that over time the $R$ drifts upwards towards the key-worker compartment $R$. To avoid artefacts, models without separate locked down and key worker compartments clearly need to include an $R$ model flexible enough to capture this expected dip and recovery in $R$. Birrell et al. (2021) did this, but the highly cited Flaxman et al. (2020) did not, instead assuming that $R$ was constant during lockdown: the serious artefacts that this induced are discussed in section 5.3.
@@ 5 Lockdowns
$f(t)$ is the log fatal incidence rate at day $t$ and $\pi(d)$ is the probability of an infection to death time interval of $d$ days. $D_{i}$ is the maximum lag from infection to death considered. To promote statistical stability, at the start of the epidemic this may be set to somewhere around 20 days, since the first deaths observed will tend to be from shorter duration disease. $D_{i}$ then grows at a day per day up to some limit (e.g. 80 days). This approach avoids estimating $f$ over a long initial period where $\exp(f)$ is essentially zero. The model can also be multiplied by a second log cyclic smooth term, to deal with the slight weekly cycle in deaths seen in some countries. $y_{i}$ can be assumed to follow a negative binomial or Poisson distribution. The smooth terms in the model can be represented using cubic splines with smoothing parameters estimated by (Laplace approximate) REML. Assuming smoothness on the log scale mitigates against the possibility of smoothing artefacts driven by rapid changes in absolute incidence. Appendix B provides more detail.
@@ 6 Discussion
The detachment from external reality promoted by the online environment provides fertile ground for the post-modern idea that language in fact creates, controls or is reality, at least in the social sphere, perhaps feeding the cognate notion that mathematics, the language of science, is equivalent to or controls scientific reality. The ‘illusion of control’ that this creates (e.g. Gupta, 2001) is seductive, but unhelpful if it results in an excessive effort being devoted to modelling rather than measurement. This is not an argument against quantitative science. For example, the scientifically advised central UK government did not fall for the deeply unscientific and innumerate belief (see Dowdle, 1998) that an endemic disease could, for the first time in history, be eliminated by physical distancing measures, if only these were sufficiently stringent and prolonged (‘zero Covid’). Sticking to the quantitative science, in this respect, avoided the even greater collateral damage of harsher more prolonged measures.
@@ Appendix A Iterating life table demography
was estimated from 2017-19 data, where $f_{1}$ is a cyclic smooth function and $f_{2}$ is a centred slowly varying smooth function, while $\sigma$ and $\nu$ are parameter to be estimated. Then $d_{w}=\hat{f}_{1}(w)/\sum_{w=1}^{52}\hat{f}_{1}(w)$ defines the multiplier of average weekly mortality required to account for seasonal variation.
@@
When iterated for 3 years from the estimated population by age at the start of 2017 this approach slightly underestimates actual deaths by just under 50 out of 1.8 million ($<0.0026\%$). This slight underestimation will lead to a slight overestimation in excess deaths (somewhere around $0.05\%$). Code and data used are provided in the supplementary material.
@@ Appendix B Modelling with smooth functions
[…]
@@
where ${\bf V}_{\beta}=(\partial^{2}l/{\partial{\bm{\beta}}\partial{\bm{\beta}}^{\sf T}}+\lambda{\bf S})^{-1}$. Writing $\pi_{G}({\bm{\beta}}|{\bf y})$ for this Gaussian approximation to the posterior and $\pi({\bm{\beta}})$ for the smoothing prior, the marginal likelihood is approximately $\exp\{l(\hat{\bm{\beta}})\}\pi(\hat{\bm{\beta}})/\pi_{G}(\hat{\bm{\beta}}|{\bf y})$ (Laplace approximation) which can be maximized to estimate $\lambda$. This approach is equally applicable when there are several smoothing parameters and a penalty of the form ${\bm{\beta}}^{\sf T}{\bf S}_{\lambda}{\bm{\beta}}$ where ${\bf S}_{\lambda}=\sum\lambda_{j}{\bf S}_{j}$.

### v3

@@ 4.1 Person-to-person variability in SEIR models
where $I(t)=\int i(\alpha,t)d\alpha$. On integrating the first ODE we have $s(\alpha,t)=s(\alpha,0)\exp(-\alpha q_{t})$ where $q_{t}=\int_{0}^{t}I(t^{\prime})dt^{\prime}$. Since $q_{t}$ is monotonic in $t$, it is immediately clear how the epidemic progresses faster in subpopulations with higher $\alpha$. In itself this observation suggests that great care is needed in extrapolating to the whole population from those who become sick first.
@@
$\frac{{\rm d}S}{{\rm d}t}=-\int\alpha s(\alpha,t)d\alpha I(t)=-M^{\prime}(-q_{t})I(t)=-M^{\prime}\{M^{-1}(S_{t})\}I(t)$
@@
$\frac{{\rm d}s(\alpha,t)}{{\rm d}t}=-\int\alpha\alpha^{\prime}s(\alpha,t)i(\alpha^{\prime},t)d\alpha^{\prime}=-\alpha\bar{\alpha}^{\prime}_{t}s(\alpha,t)I(t)$
where $\bar{\alpha}^{\prime}_{t}=\int\alpha^{\prime}i(\alpha^{\prime},t)/I(t)d\alpha^{\prime}$. Analytical progress now requires the approximation that the infectious state is short enough that the distribution of $\alpha$ in the infectious stage at $t$ is proportional to the distribution in those first becoming infected at $t$. That is $i(\alpha,t)=\alpha s(\alpha,t)$, so that $\bar{\alpha}^{\prime}_{t}=\int\alpha^{2}s(\alpha,t)d\alpha/\int\alpha s(\alpha,t)d\alpha$. If we now redefine $q_{t}=\int_{0}^{t}\bar{\alpha}_{t^{\prime}}^{\prime}I(t^{\prime})d\alpha^{\prime}$, then the maths follows through similarly to the variable susceptibility case, with the addition that $\bar{\alpha}^{\prime}_{t}=M^{\prime\prime}(-q_{t})/M^{\prime}(-q_{t})$, so that we end up with
@@ 4.2 Other major modelling omissions
One such consequence concerns the modelling of $R$ after lockdown. $R$ measures the average number of new infections caused by each existing infection. Crucially it is the population of infections that is being averaged over, not the population of people. Immediately after lockdown $R$ is depressed in the locked-down population, increasingly so with time as the infectious run out of household members to infect. Meanwhile the key-worker population maintains a higher $R$. Since initially most infections are in the locked-down compartment the whole population $R$ initially tracks what is happening in that compartment, but over time an ever increasing proportion of infections is in the key-worker population, so that over time $R$ drifts upwards towards the key-worker compartment $R$. To avoid artefacts, models without separate locked down and key worker compartments clearly need to include an $R$ model flexible enough to capture this expected dip and recovery in $R$. Birrell et al. (2021) did this, but the highly cited Flaxman et al. (2020) did not, instead assuming that $R$ was constant during lockdown: the serious artefacts that this induced are discussed in section 5.3.
@@ 5 Lockdowns
$f(t)$ is the smooth log fatal incidence rate at day $t$ and $\pi(d)$ is the probability of an infection to death time interval of $d$ days. $D_{i}$ is the maximum lag from infection to death considered. To promote statistical stability, at the start of the epidemic this may be set to somewhere around 20 days, since the first deaths observed will tend to be from shorter duration disease. $D_{i}$ then grows at a day per day up to some limit (e.g. 80 days). This approach avoids estimating $f$ over a long initial period where $\exp(f)$ is essentially zero. The model can also be multiplied by a second log cyclic smooth term, to deal with the slight weekly cycle in deaths seen in some countries. $y_{i}$ can be assumed to follow a negative binomial or Poisson distribution. The smooth terms in the model can be represented using cubic splines with smoothing parameters estimated by (Laplace approximate) REML. Assuming smoothness on the log scale mitigates against the possibility of smoothing artefacts driven by rapid changes in absolute incidence. Appendix B provides more detail.
@@ 6 Discussion
The detachment from external objective reality promoted by the online environment provides fertile ground for the post-modern idea that language in fact creates, controls or is reality, at least in the social sphere, perhaps feeding the cognate notion that mathematics, the language of science, is equivalent to or controls scientific reality. The ‘illusion of control’ that this creates (e.g. Gupta, 2001) is seductive, but unhelpful if it results in an excessive effort being devoted to modelling rather than measurement. This is not an argument against quantitative science. For example, the scientifically advised central UK government did not fall for the deeply unscientific and innumerate belief (see Dowdle, 1998) that an endemic disease could, for the first time in history, be eliminated by physical distancing measures, if only these were sufficiently stringent and prolonged (‘zero Covid’). Sticking to the quantitative science, in this respect, avoided the even greater collateral damage of harsher more prolonged measures.
@@ Appendix A Iterating life table demography
was estimated from 2017-19 data, where $f_{1}$ is a cyclic smooth function and $f_{2}$ is a centred slowly varying smooth function, while $\sigma$ and $\nu$ are parameters to be estimated. Then $d_{w}=\hat{f}_{1}(w)/\sum_{w=1}^{52}\hat{f}_{1}(w)$ defines the multiplier of average weekly mortality required to account for seasonal variation.
@@
When iterated for 3 years from the estimated population by age at the start of 2017 this approach slightly underestimates actual deaths by just under 50 out of 1.8 million ($0.0026\%$). This slight underestimation will lead to a slight overestimation in excess deaths (somewhere around $0.05\%$). Code and data used are provided in the supplementary material.
@@ Appendix B Modelling with smooth functions
Both the death deconvolution models and the replications of Flaxman et al. (2020) and Knock et al. (2020), covered in section 5, are statistical models in which smooth functions of time are to be estimated alongside other parameters. A basis expansion is employed for the smooth function, $f(t)=\sum_{k=1}^{K}\beta_{k}b_{k}(t)$ where $\beta_{k}$ is an unknown coefficient targeted by statistical inference and $b_{k}(t)$ a known basis function chosen for good approximation theoretic properties. A cubic spline basis is convenient. Associated with $f(t)$ is a smoothing penalty, such as $\lambda\int f^{\prime\prime}(t)^{2}dt=\lambda{\bm{\beta}}^{\sf T}{\bf S}{\bm{\beta}}$ ($\bf S$ known), which can be used to penalize complexity of $f$ during inference, tuneably via the smoothing parameter $\lambda$. In a Bayesian setting it is natural to view such a penalty as being induced by an improper Gaussian smoothing prior ${\bm{\beta}}\sim N({\bf 0},{\bf S}^{-}/\lambda)$.
@@
where ${\bf V}_{\beta}=(-\partial^{2}l/{\partial{\bm{\beta}}\partial{\bm{\beta}}^{\sf T}}+\lambda{\bf S})^{-1}$. Writing $\pi_{G}({\bm{\beta}}|{\bf y})$ for this Gaussian approximation to the posterior and $\pi({\bm{\beta}})$ for the smoothing prior, the marginal likelihood is approximately $\exp\{l(\hat{\bm{\beta}})\}\pi(\hat{\bm{\beta}})/\pi_{G}(\hat{\bm{\beta}}|{\bf y})$ (Laplace approximation) which can be maximized to estimate $\lambda$. This approach is equally applicable when there are several smoothing parameters and a penalty of the form ${\bm{\beta}}^{\sf T}{\bf S}_{\lambda}{\bm{\beta}}$ where ${\bf S}_{\lambda}=\sum\lambda_{j}{\bf S}_{j}$.